OpenStudy (anonymous):

Calculate the percent composition of Mg(NO3)2.

OpenStudy (abb0t):

2 calculate the percent composition, determine the molar mass of the compound In this case, the molar mass \[24.31 + 2 \times [ 14.01 + (3 \times 16.00)] = ?????\] then divide the molar mass by the individual mass of each element. including the number of moles present in the elements.